# Towards User-Focused Research in Training Data Attribution for Human-Centered Explainable AI

### v1

@@ 4.3.3. Axis 3: Number of training samples
The third axis corresponds to the number of training samples participants need to make sense of a model error through the training data. We identify two main options: A single sample and a group of samples (z-axis of Figure 3). We found that model developers usually do not inspect just one training data sample but think more globally as they usually refer to training data and samples in plural form (#10). This aspect is addressed in TDA research by works studying group attribution (Koh et al., 2019; Basu et al., 2020; Park et al., 2023). By understanding preferences in this axis, we aim to identify whether individual attribution like in traditional TDA or rather training data groups need to be in focus of TDA research.
4.4. Survey study
Building on the findings of the interview study, we aim to study the needs of model developers and understand what kind of information is useful, addressing RQ2 – What do users need from TDA explanations? We design an online survey that explores the design space of TDA explanations (see Figure 3) using scenario-based design (Carroll, 1995). We piloted the survey to ensure clear phrasing and compatibility across devices. Data was collected from June – August 2024.
@@ 4.4.1. Participants
We recruit participants with prior experience in developing machine learning models to closely align with potential real users. To recruit expert participants, we used snowball sampling alongside social media ads on X and LinkedIn, reaching out to contacts and asking them to share the study link. We expect thematic saturation in the qualitative analysis and meaningful statistics in the quantitative analysis at around 30 replies. Of 34 responses, we excluded three for low-quality answers (short and non-sensical), leaving 31 participants (nine female). 18 participants work in industry while 14 work in academia, with most having 1-5 years of machine learning experience. Four do not work with deep models. Participants come from diverse fields, including biology, health, logistics, and more. While the data types used by participants are diverse too, the majority of participants work with image or tabular data, as used in our study (see Figure 4). The participants were compensated with 25€ via Tremendous111https://www.tremendous.com/.
This figure shows five histograms about the answer distributions for the preliminary demographic questions of the survey study. From the top right corner to the bottom left corner, the histograms show the gender distribution (22 male vs. 9 female), whether people work in industry or academia (13 academia, 17 industry, 1 in both), years of experience distribution (1 ¡1 year, 15 1-3 years, 12 3-5 years, 3 ¿5years), the distribution of data types that participants work with (19 Image, 16 tabular, 14 time series, 15 text, 8 video, 1 biological, 2 graphs, 1 audio and 1 robot data), the distribution of experiences with deep learning models (27 yes vs. 4 no).
@@ 4.4.2. Scenario-based design
The main part of the study is constructed with scenario-based design (Carroll, 1995) to base the analysis on a tangible usage scenario. We create two scenarios of model debugging in an interactive mock-up of a model development suite (see Figure 5). The choice of data types (image and tabular data) is based on the dominant modalities of interview participants. The study puts participants in two imaginary scenarios where they are model developers in a company that builds (1) a bird classification app, and (2) a credit scoring app. The company recently acquired a data-centric tool to help model developers debug their models by understanding errors and identifying training data relevant to those errors. The tool is customisable to adapt to the developer’s preferences. The customisation choices, based on the TDA design space (see Figure 3) are:
@@
In both scenarios, participants are presented with a specific misclassification as the model error (left side of Figure 5). This error is randomised across participants to prevent bias from a specific sample. Participants are asked to choose which type of training data-related information (right side of Figure 5) is most useful to understand the reasons for the error. This way, we aim to elicit participants’ intuitions about the most actionable information for debugging the model error. Participants are encouraged to explore different settings and submit their desired customisation to complete the task. The data for these scenarios was generated using real machine learning models and openly available datasets. We detail the data generation process in Appendix C.
The figure shows a screenshot of the bird classification scenario interface of the survey study. The page is divided into two parts: On the left is the interface with the header “Customize your own development interface”, and on the right are the customization controls as an expanded menu. The controls have the instructions: “Choose a setting below and click Apply to see the effects.” Below are three drop down lists, one for Action defining relevance, one for Metric of interest and the last one for the number of training samples. The settings are set to “Removal”, “Loss” and “1” respectively. Below the dropdown lists is a grey button with the word Apply on it. The interface itself is divided into two panels as well: on the left, indicated with a “1” is a panel with the title “Test error analysis”. Below, the test accuracy is given as well as a description of the Error (“Wren is misclassified as Sparrow.”). Underneath the error description is the image of the wren that is misclassified with information about the prediction, loss, predictive probability and ground truth probability. The second panel, indicated with a “2” is titled “Training data analysis”. It shows both the individual effect (how does the model behaviour on the erroneous test sample change?) of the setting chosen in the customisation menu as well as the global effect (how does the overall model behaviour change?). Below the effect is a table with the most relevant training samples according to the chosen setting, with their ground truth and predicted labels. The dynamic information is highlighted by an orange box with arrows pointing from the customisation menu to the dynamic panel.
4.4.3. Survey structure
@@
The survey structure is shown in Figure 6. Participants are first briefed on the study’s objective and data processing. After providing consent and confirming they meet inclusion criteria, they are asked preliminary demographic questions for context to the later analysis. In the remaining study, we aim to extract the participant’s preferences through a usage scenario (§ 4.4.2). The participants are presented with the scenario description and prompted with “Imagine you are a model developer in a company […]” to immerse them in the use case. Then, we ask the participants follow-up questions to understand the reasoning behind their choices through free-text questions. These free-text questions are about understanding the participants’ intuitions of why the model may have made an error, and how they would fix it. After completing one use case, the participants are presented with the other use case, where the order is randomised across users to prevent an order bias.
The figure shows a flowchart of the survey. The survey starts with a question about prior model developing experience to assess whether inclusion criteria are met. Then, the next state of preliminary questions follow to record context to the data. This includes demographic questions and questions about prior model development experiences. The third state is the tutorial for the interactive application, after which the quantitative data collection is started with either the bird classification (image) use case or the credit scoring (tabular) use case. In this stage, participants can explore the interface and are asked to submit their final choices. The next stage is the qualitative data collection, where participants are asked to explain their choices and proposed actions in free text. Then, an arrow indicates that the other use case will be presented before the end of the survey.
@@ 4.4.4. Data analysis
We analyse the survey study both quantitatively and qualitatively. For the quantitative analysis, we aim to understand whether a specific notion of TDA is preferred across participants because previous TDA studies have been centred around a single notion of TDA (§ 3). Since the traditional TDA setting is already a part of the design space, studying preferences gives an answer to whether participants really preferred this particular scenario. We further assess if participants prefer a specific scenario —- if not, it is a strong signal that the research should be diversified into multiple possible scenarios serving different user needs.
To this end, we compute the distribution of the participants’ choices (action, metric and number of samples). Furthermore, we test whether the collected data indicates a statistically significant preference for a specific setting. We perform a test as the data is categorical (Pearson, 1900). For the number of training samples axis, we perform Fisher’s exact test as the choice frequencies are low (Edwards, 2005). Specifically, our hypotheses are:
@@
The qualitative analysis aims to understand the reasons behind participant choices and get deeper insight into user preferences. The free-text responses in the follow-up questions are analysed using inductive thematic coding (Braun and Clarke, 2012) by two coders. Per question, we analyse the common themes among participants to extract the reasons behind user preferences and needs. The analysis consists of several coding workshops and is depicted in Figure 7. We measure interrater agreement by the percentage of overlapping themes. The final agreement is 97.1%.
The figure shows the progression of the thematic analysis of the survey data in a timeline-like schematic consisting of 8 parts: The initial coding workshop, where the data of five participants was analysed, resulted in 21 themes. The next stage is independent coding of eight more participants, followed by the first intermediate coding workshop that resulted in 29 themes with an IRR of 75.8%. This is repeated and the next intermediate coding workshop resulted in 31 themes with an IRR of 84.3%. Then, four more participants are independently analysed before the final coding workshop that resulted in 32 themes and and IRR of 97.1%. Lastly, the stage of case-by-case discussions of themes indicates that single participants who answered the survey after the main analysis are discussed on a case-by-case basis.
4.5. Survey findings
We conduct the survey study to answer RQ2 – What do users need from TDA explanations? The survey data analysis indicates a clear connection between users’ intuition of what causes a model error and their TDA preferences. We elaborate further in the following.
@@ 4.5.1. Quantitative results
The quantitative analysis builds on the customisation choices of the debugging interface (see Quantitative data collection in Figure 6). Table 3 shows the frequency of selections and the associated p-values. We test if there exists a setting of action, metric and number of training samples which is consistently chosen and particularly useful to model developers. We are especially interested in whether there is a clear preference for removal, loss and one training sample as these correspond to the traditional notion of TDA introduced in § 3. For the action axis, there is no statistically significant preference for any choice and the distribution is rather uniform across choices. While the results may show a higher frequency of participants choosing loss as a metric, we observe no statistically significant preference. Hence, we fail to reject for the action and metric axes and conclude that there is no dominant preference for a particular TDA setting. For the number of samples axis, however, we find a statistically significant preference for a group of 10 samples.
The entropy analysis supports this conclusion (see Table 4). As a measure of diversity, normalised entropy scores range from 0.588 to 0.997, showing a highly diverse (¿0.5) distribution of choices. This is especially true for the Action defining relevance and Metric axes with high scores (¿0.8). There is less diversity in the number of training samples but a clear preference for choosing groups of samples, confirming Ilyas et al.’s (Ilyas et al., 2022) intuitions for studying group attribution.
@@
| Removal | 33.33% | 1 | 41.97% | 1 | 37.70% | 1 |
| Label Change | 30.00% | 1 | 16.13% | 1 | 22.95% | 1 |
| Upweighting | 36.67% | 1 | 41.94% | 1 | 39.34% | 1 |
@@
| Loss | 60.00% | 1 | 35.48% | 1 | 47.54% | 1 |
| Pred. cls. prob. | 10.00% | 1 | 41.94% | 1 | 26.23% | 1 |
| GT class prob. | 30.00% | 1 | 22.58% | 1 | 26.23% | 1 |
@@ 4.5.2. Qualitative results
The figure shows the themes found in the coding analysis of the survey study, organised into four theme areas. The first theme area is called “Action depends on user’s hypotheses” and includes the themes on different hypotheses that participants have about the root cause of the model error (mislabelled data, incomplete training, spurious correlations, similar class features, low dataset diversity, class overlap, dataset imbalance and small dataset size) as well as themes related to the proposed bugfixing actions users proposed (Remove noisy data, changing label to correct, fix labelling pipeline for future data, training process engineering, model architecture engineering, feature engineering, upweight data to focus on specific features, feature attribution inspection, manual inspection of datapoints, collect specific new data, synthesize/augment data, resample/reweigh data and collect any new data). The second theme area is called “Individual metric preferences” and includes themes on preferred metrics (probabilities are intuitive and show the decision boundary, ground truth probability is the target, loss is most informative in the multi-class case) and one theme related to preferred actions (Label change together with a probability metric show the decision boundary). The third theme area is titled “Group attribution preferences” with themes Group for overview of the error, Balance betw. variety and conciseness. The last theme area is titled “Reliability of explanations” and includes themes related to the preferred number of training samples (group size enough for reliable attribution, small group to prevent reasoning change), preferred metric (Ground truth probability is the target) and preferred action (Removal shows causal effect).
The inductive thematic coding analysis of the free-text answers resulted in four thematic areas and theme types covering 30 themes (see Figure 8): (1) Action depends on the user’s hypothesis, (2) Individual metric preferences, (3) Group attribution preferences, and (4) Reliability of explanations. For the theme types, we find themes that correspond to the user’s hypothesis about the error and observe that these are often coded together with a corresponding proposed action. However, preferred metrics and preferred number of training samples do not correspond to the user’s hypothesis. Additionally, we identify four themes that do not correspond to any category but give interesting insights and context to the analysis. The full description of each theme is in Appendix D.
Action depends on user’s hypotheses.
@@
The thematic analysis complements our quantitative findings: User preferences in model debugging are highly individual. The analysis provides a reason: Individual preferences are often linked to the user’s understanding and intuitions, in other words, their hypothesis about an error. The user’s hypothesis determines their course of action. Hypotheses are highly individual because users formulate them based on their expertise and past experiences. For example, we identified the theme of low dataset diversity as a hypothesis participants have about a model error: “[Images] […] in the training set are not diverse enough”, “[One class] was not representative enough”. Corresponding to this hypothesis, participants often suggested adding new data or applying data augmentations to diversify the training set, which we identified as collect specific new data and synthesise/augment data themes.
Individual metric preferences.
@@
For individual metric preferences, we identify a link with the task at hand. In the case of binary classification, participants showed a preference for probability metrics, as it is easier to understand than the loss and allows an inspection of the decision boundary (“The loss is not very intuitive”). In multi-class classification, more participants preferred the loss as it considers all possible classes (“[The loss] contains information on all classes”). However, as the quantitative analysis shows, these differences in preference are not statistically significant. Therefore, the metric preferences are likely individual to the user and their habits.
Group attribution preferences.
@@
The quantitative study shows a clear preference for attribution to groups of training samples as opposed to individual training samples. The qualitative analysis finds that this is due to the user’s need for having a good overview of the model error. Inspecting a group of training samples provides a better overview than a single sample, and allows the user to understand whether their hypothesis about the error is valid or not (Group for overview of the error). A second theme highlights another angle: While variety in the group is desired for a good overview of the error (e.g. group consisting of samples with diverse features), the group should be concise enough to “keep the analysis manageable” (Balance betw. variety and conciseness).
Reliability of explanations.
Our analysis finds that some themes are related to the overall concept of explanation reliability, where particularly the choice of how many training samples to inspect is associated with reliability. Participants stated that group sizes should be large enough for the attribution scores to be reliable. At the same time, he group size should be small enough to ensure reliable attribution scores, as changing the dataset could change the model’s reasoning (Small group to prevent reasoning change). We interpret this need for reliability to reflect in the motivation to uncover causal effects (“By removing […] I can determine if [the training samples] were causing confusion within the model”).
Unintended themes.
We identified four unintended themes outside the study design that provide additional context. We noticed that participants focused either on the model error (Focus on model error) or the final accuracy (Focus on final accuracy). Therefore, some of the participants explored all choice options in the survey and submitted the setting with the largest performance gain. This indicates that the end goal of performance improvement in the task is highly valued. Some participants mentioned that the study interface is unfamiliar (Unfamiliar study interface), which can be confusing. We considered this point in the analysis through discussions at the coding workshops. However, it also indicates that participants were unfamiliar with TDA explanations overall as they are not used in practice yet. The last theme highlights the importance of grouping erroneous test samples. Participants noted that a single erroneous test sample may represent a broader issue, and understanding the link between training data and groups of test samples with similar errors could enhance debugging.

### v2

@@ 4.3.3. Axis 3: Number of training samples
The third axis corresponds to the number of training samples participants need to make sense of a model error through the training data. We identify two main options: A single sample and a group of samples (z-axis of Figure 4). This axis is inspired by the explicit answer of I5 (“I think [the most relevant data] would be like a collection of [samples].”) and the implicit indication of talking about multiple samples in the participants’ answers, recognisable through the plural form use of “data”, “samples” and “artefacts” (Theme: Intuitions on data explanations Code: Group of samples).
We found that model developers usually do not inspect just one training data sample but think more globally in groups of relevant data. This aspect is addressed in TDA research by works studying group attribution (Koh et al., 2019; Basu et al., 2020; Park et al., 2023; Hu et al., 2024). By understanding preferences in this axis, we aim to identify whether individual attribution, like in traditional TDA, or rather training data groups, needs to be in focus of TDA research.
4.4. Scenario-based Interactive User Study
Building on the findings of the interview study, we designed an interactive online study that explores the design space of TDA explanations (see Figure 4) using scenario-based design (Carroll, 1995). We piloted the study interface to ensure clear phrasing and compatibility across devices. Data was collected from June – August 2024.
@@ 4.4.1. Participants
We recruited participants with prior experience in developing machine learning models to closely align with potential real users. Contrary to the interview study, we did not constrain our participants’ work to be in a high-risk application domain for improved reach. To recruit expert participants, we used snowball sampling alongside social media ads on X and LinkedIn, reaching out to personal contacts and asking them to share the study link. We aimed for around 30 replies to provide sufficient diversity in the answers for the qualitative analysis while also being indicative of measurable effects for a quantitative analysis.444We conducted a test of statistical significance. A power analysis at significance level , power and large effect size yields a sample size of 31. Of 34 responses, we excluded three for low-quality answers (short and non-sensical), leaving 31 participants (nine female). 18 participants work in industry while 14 work in academia, with most having 1-5 years of machine learning experience. Four do not work with deep models. Participants come from diverse fields, including biology, health, logistics, and more. While the data types used by participants are diverse too, the majority of participants work with image or tabular data, as used in our study (see Figure 5). The participants were compensated with 25€ (=$27.05) via Tremendous555https://www.tremendous.com/.
@@ 4.4.2. Scenario-based design
The main part of the study is constructed with scenario-based design (Carroll, 1995) to base the analysis on a tangible usage scenario. We created two scenarios of model debugging in an interactive mock-up of a model development suite (see Figure 6). The choice of data types (image and tabular data) was based on the dominant modalities of interview participants. The study put participants in two imaginary scenarios where they are model developers in a company that builds (1) a bird classification app, and (2) a credit scoring app. The company recently acquired a data-centric tool to help model developers debug their models by understanding errors and identifying training data relevant to those errors. The tool is customisable to adapt to the developer’s preferences. The customisation choices, based on the TDA design space (see Figure 4) were:
@@
In both scenarios, participants were presented with a specific misclassification as the model error (left side of Figure 6). This error was randomised across participants to prevent bias from a specific sample. Participants were asked to choose which type of training data-related information (right side of Figure 6) is most useful to understand the reasons for the error. This way, we aimed to elicit participants’ intuitions about the most actionable information for debugging the model error. Participants were encouraged to explore different settings and submit their desired customisation to complete the task. The data for these scenarios was generated using real machine learning models and openly available datasets (Wah et al., 2011; Hofmann, 1994). We detail the data generation process in Appendix B.
4.4.3. Study structure
@@
The structure of the scenario-based interactive study is shown in Figure 7. Participants were first briefed on the study’s objective and data processing. After providing consent and confirming they meet inclusion criteria, they were asked preliminary demographic questions for context to the later analysis. In the remaining study, we aimed to extract the participant’s preferences through a usage scenario (§ 4.4.2). The participants were presented with the scenario description that details their objective of understanding possible reasons for a presented error. They were prompted with “Imagine you are a model developer in a company […]” to immerse them in the use case. Then, we asked the participants follow-up questions to understand the reasoning behind their choices through free-text questions (e.g.“Why did you choose to look at the {chosen_metric} metric?”). These free-text questions were about understanding the participants’ intuitions of why the model may have made an error (“In your opinion, why did the model fail at the above classification?”), and how they would fix it (“What would you do to fix this error?”). After completing one use case, the participants were presented with the other use case, where the order is randomised across users to prevent an order bias.
@@ 4.4.4. Data analysis
We analyse the scenario-based interactive study both quantitatively and qualitatively. The quantitative analysis aims to examine the diversity of participant preferences in TDA explanations. Specifically, we investigated whether the predominant form of explanation in TDA, i.e. the counterfactual effect of removing a sample (§ 3), aligns with the explanatory needs of model developers. If there is alignment, TDA research focuses on the right problem, and methods produced in this field are valuable to model developers. However, if this is not the case, it is a strong signal that the research should be diversified into multiple possible scenarios serving different user needs.
To this end, we computed the distribution of the participants’ choices (action, metric and number of samples). We tested whether the collected data indicates a statistically significant preference for a specific setting. We performed a test per dimension of the design space (action defining relevance, metric, number of training samples) as the data is categorical (Pearson, 1900). We considered two categories for the number of training samples, the first one being a single sample and the second one being groups larger than one as defined in the design space (Figure 4). Specifically, our hypotheses are:
@@
The qualitative analysis aims to understand the reasons behind participant choices and get deeper insight into user preferences. The free-text responses in the follow-up questions were analysed using inductive thematic coding (Braun and Clarke, 2012) by two coders. Per question, we analysed the common themes among participants to extract the reasons behind user preferences and needs. The analysis consisted of several coding workshops and is depicted in Figure 8. We measure interrater agreement in Krippendorff’s (Krippendorff, 2011). The final agreement is .
4.5. Study findings
We conducted the interactive study to study model developers’ TDA needs through a scenario. The data analysis indicates a clear connection between developers’ intuition of what causes a model error and their TDA needs. We elaborate further in the following.
@@ 4.5.1. Quantitative results
The quantitative analysis builds on the customisation choices of the debugging interface (see Quantitative data collection in Figure 7). Table 2 shows the frequency of selections and the associated p-values. We test whether a specific action, metric, and number of training samples are consistently chosen and particularly useful to model developers. We are especially interested in a clear preference for removal, loss, and one training sample as these correspond to the predominant notion of TDA introduced in § 3. The analysis yields the following results: For the action axis, there is no statistically significant preference for any choice, and the distribution is rather uniform across choices. While the results may show a higher frequency of participants choosing loss as a metric, we observe no statistically significant preference. Hence, we fail to reject for the action and metric axes and conclude that there is no dominant preference for a particular TDA setting. Instead, developers have individual preferences. For the number of samples axis, however, we observe that all of our participants chose ¿1 samples across both scenarios. More specifically, we find a statistically significant preference for a group of 10 samples, indicating a clear preference for training data groups instead of individual samples.
The entropy analysis further supports this conclusion (see Table 3). As a measure of diversity, normalised entropy scores range from 0.588 to 0.997, showing a highly diverse (¿0.5) distribution of choices. This is especially true for the action and metric axes with particularly high scores (¿0.8). There is less diversity in the number of training samples but a clear preference for choosing groups of samples, confirming Ilyas et al.’s (Ilyas et al., 2022) intuitions for studying group attribution.
@@
| Removal | 33.3% | 1 | 42.0% | 1 | 37.7% | 1 |
| Label Change | 30.0% | 1 | 16.1% | 1 | 23.0% | 1 |
| Upweighting | 36.7% | 1 | 41.9% | 1 | 39.3% | 1 |
@@
| Loss | 60.0% | 1 | 35.5% | 1 | 47.5% | 1 |
| Pred. cls. prob. | 10.0% | 1 | 41.9% | 1 | 26.2% | 1 |
| GT class prob. | 30.0% | 1 | 22.6% | 1 | 26.2% | 1 |
@@ 4.5.2. Qualitative results
The inductive thematic coding analysis of the free-text answers resulted in eight themes that address the reasoning behind participants’ choices for action, metric, and number of samples preferences (see Figure 9). We elaborate on the themes below.
Theme: Error hypotheses informs action
@@
The thematic analysis complements our quantitative findings: Developer preferences in model debugging are highly individual. The analysis provides a reason in line with the sensemaking framework introduced in Cabrera et al. (Cabrera et al., 2023): Individual preferences for proposed actions are often linked to the developer’s understanding and intuitions. Needs are highly individual because developers may have different mental models based on their expertise and past experiences. In other words, their mental models of the model’s learned behaviour affect their hypothesis about an error and the chosen action. The developer’s hypothesis determines their explanatory needs: An explanation needs to enable the developer to reject or confirm their hypothesis of why a model error occurs so that they can make informed decisions and actions to resolve the error. For example, participants who hypothesise the cause of an error to be data-related also propose data-related actions: “[Images] […] in the training set are not diverse enough”, “[One class] was not representative enough”. Corresponding to this hypothesis, participants often suggested adding new data or applying data augmentations to diversify the training set. If the hypothesis is linked to the model setup, the proposed action corresponds to the model.
Themes about type of actions
In terms of types of actions, we identified three main themes: Data-based actions, model-based actions, and inspection actions. In particular, the data-based actions are relevant to studying TDA needs (Theme: Data-based actions). In line with the interview study findings, we found that there are various data-based actions that participants propose: From removing malicious data to balancing the dataset through reweighting or adding new data. Additionally, we found that participants propose synthetically creating new data to add diverse data or fill up undersampled parts of the dataset (“address class imbalance by synthesizing data”). This shows that the actions participants propose are various and our collection is not likely to be exhaustive. Hence, useful and actionable TDA needs to facilitate a variety of actions that help developers verify their hypotheses about the model error. Model-based actions address changes in model architecture or hyperparameter tuning and are connected to participant hypotheses about incomplete training (“Apply regularization methods like dropout or batch normalization to improve generalization and reduce overfitting.”) (Theme: Model-based actions). Inspection actions are linked to first understanding a suspected bias of the model to get a clearer hypothesis about the error (Theme: Inspection actions). We found feature attribution explanation approaches (“Check which features the model considers important using techniques like SHAP […].”) and manual inspection-based approaches (“Typically, I would focus on the specific test sample and try to understand why the model is making a mistake by analyzing the feature values and their relationships and checking for outliers or anomalies.”). TDA explanations were not mentioned, but could ideally support these inspection activities.
Themes about metrics
@@
We identified a link between the preferred metric and the task at hand. In the case of binary classification, participants showed a preference for probability metrics, as it is easier to understand than the loss and allows an inspection of the decision boundary (“The loss is not very intuitive”, Theme: Metric informative of decision boundary). In multi-class classification, more participants preferred the loss as it considers all possible classes (“[The loss] contains information on all classes”, Theme: Metric informative of training objective). However, as the quantitative analysis shows, these differences in preference are not statistically significant. Therefore, the metric preferences and actionability are likely individual to the user and their habits.
Themes about number of training samples
@@
The quantitative study showed a clear preference for attribution to groups of training samples as opposed to individual training samples. The qualitative analysis found that this is due to the developer’s need for having a good overview of the model error. Inspecting a group of training samples provides a better overview than a single sample, and allows the user to understand whether their hypothesis about the error is valid or not (Theme: Group for global effect). A second theme highlights another angle: While variety in the group is desired for a good overview of the error (e.g. group consisting of samples with diverse features), the group should be concise enough to “keep the analysis manageable” and group size should be small enough to ensure reliable attribution scores, as changing the dataset could change the model’s reasoning (Theme: Group precision).
In conclusion, the analysis of the scenario-based interactive user study deepens our understanding of model developers’ needs for data-related explanations. We identify why there are diverse information needs across different developers: Since each developer may have a different mental model of their ML system, the information that is actionable to them is diverse as well and results in highly individual needs.